K1 CHOI Intro. & 02 FIN STMTS Saved Help Save & Exit Submit Check my work Avas SpinBall Corp. lists fixed assets of $32 million on its balance sheet. The firms fixed assets have recently been appraised at $56 million. Avas SpinBall Corp.s balance sheet also lists current assets at $25 million. Current assets were appraised at $46 million Current liabilities book and market values stand at $7 million and the firms book and market values of long-term debt are $28 million. Calculate the book and market values of the firms stockholders equity. Construct the book value and market value balance sheets for Avas SpinBall Corp. (Enter your answers in millions of dollars.) BOOK VALUE MARKET VALUE in millions of dollars) Book Assets Current assets Fixed assets Total Liabilities and Equity Current liabilities Long-term debt Stockholders equity Total Hint 0 $ < Prev 8 of 8 ::: Next > 27 PM

Check my work Avas SpinBall Corp. lists fixed assets of $32 million on its balance sheet. The firms fixed assets have recently been appraised at $56 million. Avas SpinBall Corp.s balance sheet also lists current assets at $25 million. Current assets were appraised at $46 million. Current liabilities book and market values stand at $7 million and the firms book and market values of long-term debt are $28 million Calculate the book and market values of the firms stockholders equity. Construct the book value and market value balance sheets for Avas SpinBall Corp. (Enter your answers in millions of dollars.) pped
Answer #1

NOTE: The appraised value is the market value whereas the listed value is the book value. Further, the solution is based on the following equation

Total Assets = Total Liabilities + Equity

( all figures in $ million ) Book Value Market Value
Assets
Current Assets 25 46
Fixed Assets 32 56
Total 57 102
Liabilities and Shareholders Equity
Current Liability 7 7
Long-Term Debt 28 28
Stockholders' Equity (57-28-7) = 22 (102-28-7) = 67
Total 57 102